OpenStudy (anonymous):

1+2+4+8+16+32+64+128+256+512=

OpenStudy (anonymous):

=1024-1=1023

OpenStudy (anonymous):

=1111111111 base 2 =1023 base10 =3FF base 12

OpenStudy (anonymous):

This is actually a geometric profession,with the common ratio 2 the sum of a geometric progression is give by \[\frac{a(r^n-1)}{(r-1)}\],where n = number of terms,a =first term and r is the common ratio. Here a = 1,n=10,so the sum is : \[\frac{1(2^{10}-1)}{2-1}=2^{10}-1=1023\]